# Umberto Collodel

### v2

@@ Abstract
This paper develops a novel method to simulate financial market reactions to European Central Bank (ECB) press conferences using Large Language Models (LLMs). We create a behavioral, agent-based simulation of 30 synthetic traders, each with distinct risk preferences, cognitive biases, and interpretive styles. These agents interpret ECB communication and forecast Euro interest rate swap levels at 3-month, 2-year, and 10-year maturities, with cross-sectional variation serving as a measure of market uncertainty. Using a comprehensive dataset of 283 ECB press conferences (1998-2025), we evaluate three prompting strategies: naive zero-shot, few-shot (enriched with historical context), and an advanced iterative ”LLM-as-a-Judge” framework. We find that even the naive approach achieves substantial correlations (roughly 0.5) between synthetic disagreement and actual post-announcement OIS volatility, particularly for medium- and long-term maturities. The LLM-as-a-Judge framework further enhances performance upon initial application, reaching correlations of almost 0.6 and identifying key drivers of volatility. Our results demonstrate that LLM-based agents can effectively capture the heterogeneous interpretive processes through which monetary policy signals propagate to financial markets, moving beyond traditional reduced-form event studies. This provides central banks with a practical tool to ex-ante evaluate communication strategies and anticipate market reactions, while offering researchers a micro-founded framework for understanding expectation formation in response to central bank communication.
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) have opened up a plethora of possibilities for economic and financial research. Due to the way these models are trained, on vast and diverse textual data, and their ability to generate contextually rich and realistic responses, LLMs can now be viewed as implicit computational representations of human behavior. Much like the economic construct of homo economicus, these models can be endowed with information, placed in hypothetical scenarios, and observed for their responses (Horton,, 2023; Kazinnik and Sinclair,, 2025). Despite their growing capabilities, the application of LLMs in finance has been relatively limited so far. [^1] A notable exception is, for example, Lopez-Lira and Tang, (2023). Yet, the financial domain may be one of the most promising areas for their use. Unlike traditional economic models that assume rational agents and efficient markets, real-world financial markets are often driven by sentiment, narrative, and behavioral biases, factors that are difficult to capture with purely mathematical models, especially in high-frequency settings. [^2] Agent-based models (ABMs) have emerged as a flexible framework to incorporate such heterogeneity (Farmer and Foley,, 2009). In the context of monetary policy, ABMs have been used to simulate expectation formation, market volatility, and policy transmission (Delli Gatti et al.,, 2011). However, most ABMs rely on stylized behavioral rules imposed by the researcher (Horton,, 2023).
A key area where LLMs can offer significant value is in modeling reactions to central bank communication, a cornerstone of modern monetary policy. It is well-established that the effectiveness of monetary policy hinges not just on the decisions themselves, but on the language used and its interpretation by economic agents. Press conferences, speeches, and official statements play a pivotal role in shaping market expectations and, consequently, asset prices (Gurkaynak et al.,, 2005; Altavilla et al.,, 2019). Previous research, has moved beyond the general direction of the market response to highlight the crucial role of response heterogeneity; in the Euro area, high disagreement following a statement transmits to financial markets through a separate channel compared to the standard monetary policy surprise channel, and can dampen the intended impact of policy tightening by lowering the term premium of the yield curve (Collodel and Kunzmann,, 2025). [^3] Bauer et al., (2021) find similar results for the US. Capturing this nuanced, belief-driven heterogeneity is a challenge for traditional models but is a natural fit for the capabilities of LLMs.
This paper introduces a novel approach to modeling market reactions to ECB communication by leveraging the capabilities of LLMs. We develop a behavioral, agent-based simulation of 30 synthetic traders through Gemini LLM, one of Google’s leading AI products. These agents respond to the texts of ECB press conferences by forecasting Euro interest rate swap levels across three key maturities, short (3m), medium (2y), and long-term (10y). The main idea is that the model can capture a range of possible interpretations of these conferences. Thus, the variation in the resulting forecasts, measured by their standard deviation, serves as a proxy for market uncertainty or disagreement. By correlating this synthetic measure with actual market outcomes, we study the ability of LLMs to replicate the heterogeneity of investors’ beliefs.
To further explore how the design of prompting influences this correlation, we implement three distinct prompting strategies. The first is a naive or “zero-shot” prompt (Radford et al.,, 2019), which serves as the baseline: each agent is assigned a simple, pre-defined level of risk aversion, a set of cognitive biases, and a unique interpretive style. The second is a “few-shot” prompt (Brown et al.,, 2020), which enriches the agent’s context by including the past three ECB press conferences along with the corresponding changes in market-based disagreement resulting from them. Finally, we introduce an LLM-as-a-Judge prompt, our most advanced strategy (Zheng et al.,, 2023). In this iterative setup, a separate LLM, the Judge, acts as a critical evaluator, systematically reviewing the analyst prompt and its resulting in-sample correlation score. The Judge then rewrites the prompt to improve alignment with market reactions. This revised prompt is fed back into the Analyst LLM, forming a feedback loop that continues over different iterations. The different prompt versions are subsequently evaluated out-of-sample to assess their robustness and decrease concerns relative to model over-fitting.
We find that LLM-based agents can effectively replicate the complex and varied reactions of investors to ECB press conferences, achieving a remarkable correlation of around 0.5 between the synthetic disagreement measure and actual market outcomes for medium- and long-term maturities, a substantial result given the inherent complexity and noise of financial markets. Contrary to expectations, a more detailed few-shot prompt does not improve this correlation; however, the LLM-as-a-Judge framework significantly enhances predictive accuracy upon its initial application, with performance degrading thereafter. These results suggest that LLM-based agent simulations may offer policymakers a useful tool for evaluating central bank communications and identifying language that could generate market disagreement prior to public release. By moving beyond traditional sentiment analysis, our approach contributes to understanding how monetary policy signals are perceived and transmitted through financial markets. The framework integrates insights from macroeconomic communication, behavioral finance, and artificial intelligence to help scholars and policymakers better understand the financial market impact of monetary policy announcements.
@@
The remainder of the paper proceeds as follows. Section 2 reviews the relevant literature on monetary policy communication, behavioral expectations, and LLM applications in economics. Section 3 outlines our methodology, including data collection of ECB press conferences and market reactions, construction of synthetic agent personas, and our three prompting strategies. Section 4 presents our empirical results, analyzing agent forecast dispersion patterns, correlation with market volatility measures, and comparative performance across prompting approaches. Section 5 probes the robustness of our results to model and prompt variation. Lastly, Section 6 concludes with policy implications for central bank communication strategies and directions for future research.
@@ 2 Literature Review
Our first contribution is to the literature on monetary policy communication, particularly the strand that identifies policy shocks using high-frequency movements in asset prices within narrow windows around policy announcements. Seminal studies for the United States show that Federal Reserve statements influence asset prices not only through policy rate changes but also through the qualitative content of the accompanying communication (Gurkaynak et al.,, 2005; Gertler and Karadi,, 2015; Hansen and McMahon,, 2016; Bauer and Swanson,, 2022). In these studies, changes in risk-free market rates measured before and after announcements serve as proxies for monetary policy shocks. In the euro area, similar evidence shows that ECB press conferences contain significant informational content beyond the policy decision itself (Altavilla et al.,, 2019). Recent work by Bauer et al., (2021) further distinguishes between the direction of policy surprises and the uncertainty surrounding the policy stance, showing that elevated uncertainty depresses equity prices, raises sovereign spreads, and increases exchange rate volatility, a pattern we also find for the Euro area (Collodel and Kunzmann,, 2025). Furthermore, Tillmann, (2020) finds that hawkish surprises are less effective in periods of heightened short- and medium-term uncertainty, as investors shift toward longer maturities in search of safety, thereby flattening the yield curve. While this literature has convincingly demonstrated that communication shocks move markets, it remains silent on how they do so. By relying on market-price changes as statistical proxies, the existing approach leaves the interpretive process of agents as a “black box.” Our work opens that box: we construct an agent-based model in which heterogeneous participants directly process the natural language of policy announcements, allowing us to trace how communication shapes expectations, generates uncertainty, and propagates into financial market dynamics.
@@
The second body of work we engage with concerns behavioral expectations. The rational expectations hypothesis, foundational to much of modern macro-finance, has been increasingly challenged by empirical evidence documenting systematic biases and bounded rationality among investors (Barberis and Thaler,, 2003; Bordalo et al.,, 2018). Standard representative-agent models struggle to reconcile such behavioral deviations with observed phenomena like excess volatility and momentum. Agent-based models (ABMs) have emerged as a flexible alternative (Farmer and Foley,, 2009), allowing for heterogeneous, adaptive agents whose interactions generate emergent macroeconomic patterns. Within monetary policy research, ABMs have been used to model expectation formation and policy transmission in bottom-up settings (Delli Gatti et al.,, 2011). Yet two key limitations remain. First, behavioral rules are typically imposed ex ante as stylized heuristics (Horton,, 2023), rather than learned or emergent. Second, conventional ABMs cannot interpret qualitative information directly, despite the fact that monetary policy signals are largely communicated in natural language. This omission bypasses an essential channel of transmission. Our framework addresses both limitations by enabling agents to process and act upon policy-relevant textual information.
@@ 3 Methodology and Data
Our methodology proceeds in three main steps: (i) construction of a structured database of ECB communication events and corresponding market reactions; (ii) development of LLM-based synthetic agents embedded in an agent-based simulation framework; and (iii) evaluation of the correspondence between synthetic measures of belief heterogeneity and actual market outcomes. We focus on ECB press conferences, which constitute the primary vehicle for conveying monetary policy decisions in the Euro area. Following the literature on high-frequency identification of monetary policy shocks (Gurkaynak et al.,, 2005; Altavilla et al.,, 2019), we define event dates as the days of regularly scheduled ECB Governing Council monetary policy meetings, limiting the sample to those that include both a written policy statement and a subsequent press conference by the ECB President. For each event, we collect the full transcript of the press conference as published on the ECB website as well as the associated OIS post-conference volatility for the short (3-months), medium (2-years), and long (10-years) term. [^5] Following the monetary policy communication literature, we use OIS rates as they reflect pure interest rate expectations without the credit risk component present in government bond yields (Gurkaynak et al.,, 2005). Our final sample comprises 283 ECB conferences between June 1998 and April 2025.
@@ 3.2 Prompting Strategies
The baseline approach, where the agent receives only the ECB press conference transcript along with its fixed attributes. This approach tests whether LLMs possess intrinsic capabilities to decode monetary policy communication without explicit calibration.
@@
Augments the baseline by including the transcripts of the three most recent ECB press conferences and the associated historical changes in market-based disagreement, providing a short memory of past communication–market interactions.This design tests whether contextual examples improve forecasting accuracy, following established few-shot learning principles. [^8] Few-shot prompting has demonstrated superior performance over zero-shot approaches across multiple economic forecasting applications, particularly when fine-tuning is computationally infeasible.
@@ 4 Results
This section evaluates the performance of our three prompting strategies in replicating market disagreement patterns following ECB press conferences. We examine the zero-shot baseline (Section 4.1), the few-shot approach with historical context (Section 4.2), and the iterative LLM-as-a-Judge framework (Section 4.3), focusing on correlations between synthetic disagreement and actual market volatility across different yield curve tenors. We conclude with a comparison against simple textual benchmarks (Section LABEL:sec:complexity) to demonstrate the value of modeling heterogeneous agent interpretations compared to less sophisticated approaches.
@@ 4.1 Zero-Shot Prompt
Note: Each data point reflects the output of an LLM model prompted in a zero-shot setting using Gemini 2.5-Flash, with a temperature parameter of 1. Forecast directions are aggregated across agents. The simulation uses a sample of 283 ECB press conference (June 1998-April 2025) transcripts and prompt 2.
@@
Note: Each data point reflects the output of an LLM model prompted in a zero-shot setting using Gemini 2.5-Flash, with a temperature parameter of 1. Dispersion is measured as the cross-sectional standard deviation of rate predictions. The simulation uses a sample of 283 ECB press conferences (June 1998-April 2025) and prompt 2. Highlighted points represent observations exceeding the 95th percentile of dispersion for each tenor.
@@
Panel 5 shows the cross-sectional standard deviation of forecasts, revealing distinct phases that align closely with monetary policy regimes and market stress episodes. Dispersion contracts systematically during periods of unambiguous policy stance, such as the post-crisis easing phase and much of the zero lower bound (ZLB) era, when forward guidance provided clear anchoring. In this period, the 3-month dispersion falls below 5 basis points, with longer tenors, instead, clustering around 10-15 basis points. Conversely, it spikes dramatically at policy turning points and during crisis episodes, touching sharp peaks during the 2008-2009 financial crisis (with the 2 and 10-year dispersion exceeding 40 basis points), the 2011-2012 sovereign debt crisis, and most recently during the 2022-2023 monetary policy tightening cycle, where disagreement reached unprecedented levels (roughly 60 basis points for the 2 and 10-year maturity). In the latter instance, disagreement arises less from directional confusion and more from uncertainty over policy speed, magnitude, and terminal endpoints. This result is consistent with the ECB 2025 Monetary Policy Strategy Review findings (ECB,, 2025). The report shows that, once forward guidance was abandoned, the sensitivity of OIS rates to macroeconomic news increased significantly, hence amplifying excess volatility. Moreover, the early phase of the 2022 hiking cycle was marked by historically large monetary policy shocks, including the unexpected 50 basis point hike in July 2022 and the largest hawkish surprise since the global financial crisis in December 2022, reflecting the challenge for the ECB of aligning communication with policy moves. By contrast, following the introduction of the three-element reaction function framework in March 2023, the subsequent phase of the tightening cycle exhibits markedly subdued volatility to policy announcements. These patterns mirror survey-based and market-implied measures of disagreement, where even directionally well-telegraphed policy paths can generate substantial heterogeneity in rate forecasts. [^17] See Gilchrist et al. (2014) and Bayer et al. (2022) for a thorough discussion on forecast disagreement and monetary policy uncertainty. [^18] At the end of 1998, volatility exhibits another peak for the 3-month tenor. In this period, the ECB avoided signaling its initial policy rate for the monetary union, deferring decisions to December. While acknowledging global risks from the Asian crisis, Russian default, and LTCM collapse, President Duisenberg reaffirmed the commitment to price stability, contrasting with the Fed’s three rate cuts between September and November. This ambiguity left markets uncertain and drove volatility in short-term instruments, mainly 1- to 3-month FRAs and DEM-based futures, since euro derivatives like Euribor futures and OIS did not yet exist (ECB,, 1999). The cross-tenor hierarchy remains remarkably consistent across all episodes, with 10-year uncertainty slightly larger than 2-year uncertainty, and both larger than the 3-month dispersion by 2-3 times, reflecting the increasing complexity of information processing and projection uncertainty at longer time horizons.
@@
Note: Correlation between the output of an LLM model prompted in a zero-shot setting using Gemini 2.5-Flash, with a temperature parameter of 1, and a market-based measure of monetary policy uncertainty from Collodel and Kunzmann, (2025). The simulation uses a sample of 283 ECB press conferences (June 1998-April 2025) and prompt 2. Rolling correlation is calculated using a 12-month window.

### v3

@@ Abstract
Traditional high-frequency identification of monetary policy communication effects operates ex-post, precluding evaluation of alternative strategies before publication. In this paper, we introduce an ex-ante framework that simulates heterogeneous market reactions to central bank communication before release. Our methodology employs Large Language Models (LLMs) to construct an agent-based simulation of 30 synthetic traders with heterogeneous risk preferences, cognitive biases, and interpretive styles. These agents process European Central Bank (ECB) press conference transcripts and forecast Euro interest rate swap rates across 3-month, 2-year, and 10-year maturities. Cross-sectional forecast dispersion provides a model-based measure of market disagreement, validated against realized overnight index swap (OIS) volatility. Analyzing 283 ECB press conferences (June 1998-April 2025), we document Spearman correlations of approximately 0.5 between simulated and realized disagreement, rising to 0.6 under iterative prompt optimization. Results prove robust across prompting strategies, are temporally stable across training and holdout samples, and fare significantly better than simple language complexity scores. For central banks, the framework provides an operational tool to anticipate communication-induced volatility before release, thus enabling ex-ante language refinement. For researchers, it offers a micro-founded alternative to reduced-form event studies, explicitly modeling the heterogeneous interpretive processes through which policy signals are transmitted to asset prices.
@@ 1 Introduction
Central banks face a fundamental communication dilemma: they cannot test how markets will interpret their statements before release. Upon release, policy language triggers immediate and irreversible market reactions. Traditional high-frequency identification methods measure these reactions ex-post, i.e., revealing what happened (Gurkaynak et al.,, 2005; Altavilla et al.,, 2019), but offering no guidance before release. This constraint is costly. Ambiguous communication can amplify market volatility, hindering financial stability and undermining policy transmission, yet policymakers lack tools to anticipate these effects during the drafting process when language remains modifiable.
This paper introduces the first ex-ante framework for simulating heterogeneous market reactions to monetary policy communication. We employ Large Language Models (LLMs) to construct an agent-based simulation of 30 synthetic traders, each endowed with distinct risk preferences, cognitive biases, and interpretive styles. These agents process European Central Bank (ECB) press conference transcripts and forecast Euro interest rate swap rates across three key maturities representing the whole yield curve spectrum (3-month, 2-year, 10-year). Cross-sectional forecast dispersion provides a model-based measure of market disagreement, which we validate against realized overnight index swap (OIS) volatility across 283 ECB communications spanning the period 1998-2025.
Our approach addresses a critical gap in the monetary policy communication literature. While extensive research documents that central bank language moves markets (Gurkaynak,, 2005; Altavilla et al.,, 2019), and that disagreement about policy stance affects transmission through distinct channels, including term premium compression (Tillmann,, 2020; Bauer et al.,, 2021; Collodel and Kunzmann,, 2025), existing methods remain fundamentally retrospective. They identify communication shocks from realized price movements, leaving the interpretive process itself as a black box. We attempt to open this box by explicitly modeling how heterogeneous agents process policy language, generating predictions of disagreement before markets react.
The framework builds on recent research demonstrating that LLMs, being trained on extensive textual corpora, can effectively simulate human interpretive behavior (Horton,, 2023; Kazinnik and Sinclair,, 2025). Unlike traditional agent-based models, which rely on stylized behavioral rules (Delli Gatti et al.,, 2011), LLM-based agents can directly process natural language, extracting meaning from the same textual signals that influence market participants. This ability is especially valuable in the context of monetary policy, where communication effects hinge on narrative and linguistic nuance rather than purely mathematical relationships. Financial markets are shaped by sentiment, biases, and interpretive style heterogeneity, features that are difficult to capture with conventional models but align naturally with the strengths of LLMs.
We implement three prompting strategies of increasing sophistication. The baseline “zero-shot” approach assigns to each agent fixed characteristics (risk aversion, cognitive biases, interpretive style) and evaluates their raw reactions to press conferences, without historical context. The “few-shot” specification enriches agents with the prior three conferences and associated market reactions, testing whether historical calibration improves prediction. Our most advanced approach employs an “LLM-as-a-Judge” framework (Zheng et al.,, 2023): a separate LLM iteratively critiques and rewrites the analyst prompt based on in-sample correlation with the market-based measure, then validates optimized prompts out-of-sample to assess overfitting.
We find that even the baseline specification achieves Spearman correlations of approximately 0.5 between simulated and realized disagreement for medium- and long-term maturities. Performance peaks at the 2-year tenor, consistent with this maturity’s elevated sensitivity to forward guidance, and rises to 0.6 under initial LLM-as-a-Judge optimization. Contrary to expectations, few-shot prompting does not improve performance. Results prove robust across model architectures. prompt choices, and are temporally stable in holdout samples.
The framework contributes to policy practice and economic research. For central banks, it provides an operational tool to anticipate communication-induced volatility before release, enabling ex-ante language refinement and counterfactual testing of alternative phrasings. For researchers, it offers a micro-founded methodology for studying expectation formation, moving beyond reduced-form event studies to explicitly model the heterogeneous belief processes through which policy signals transmit to asset prices. By demonstrating that interpretive disagreement is both quantifiable and predictable from text alone, we establish that communication strategy can be optimized systematically rather than refined only through costly trial and error.
@@
The remainder of the paper proceeds as follows. Section 2 reviews relevant literature on monetary policy communication, behavioral expectations, and LLM applications in economics. Section 3 outlines our methodology, including the construction of synthetic agents and prompting strategies. Section 4 presents empirical results on forecast dispersion patterns and correlation with market volatility. Section 5 examines robustness to model and prompt variation. Section 6 concludes with implications for central bank communication design and future research directions.
@@ 2 Literature Review
Our first contribution is to the literature on monetary policy communication, particularly the strand that identifies policy shocks using high-frequency movements in asset prices within narrow windows around policy announcements. Seminal studies for the United States show that Federal Reserve statements influence asset prices not only through policy rate changes but also through the qualitative content of the accompanying communication (Gurkaynak et al.,, 2005; Gertler and Karadi,, 2015; Hansen and McMahon,, 2016; Bauer and Swanson,, 2022). In these studies, changes in risk-free market rates measured before and after announcements serve as proxies for monetary policy shocks. [^1] For the sake of precision, the authors decompose these changes into orthogonal factors, before treating them as shocks. In the euro area, similar evidence shows that ECB press conferences contain significant informational content beyond the policy decision itself (Altavilla et al.,, 2019). Recent work by Bauer et al., (2021) further distinguishes between the direction of policy surprises and the uncertainty surrounding the policy stance, showing that elevated uncertainty depresses equity prices, raises sovereign spreads, and increases exchange rate volatility, a pattern we also find for the Euro area (Collodel and Kunzmann,, 2025). Furthermore, Tillmann, (2020) finds that hawkish surprises are less effective in periods of heightened short- and medium-term uncertainty, as investors shift toward longer maturities in search of safety, thereby flattening the yield curve. While this literature has convincingly demonstrated that communication shocks move markets, it remains silent on how they do so. By relying on market-price changes as statistical proxies, the existing approach leaves the interpretive process of agents as a “black box.” Our work opens that box: we construct an agent-based model in which heterogeneous participants directly process the natural language of policy announcements, allowing us to trace how communication shapes expectations, generates uncertainty, and propagates into financial market dynamics.
@@
The second body of work we engage with concerns behavioral expectations. The rational expectations hypothesis, foundational to much of modern macro-finance, has been increasingly challenged by empirical evidence documenting systematic biases and bounded rationality among investors (Barberis and Thaler,, 2003; Bordalo et al.,, 2018). Standard representative-agent models struggle to reconcile such behavioral deviations with observed phenomena like excess volatility and momentum. Agent-based models (ABMs) have emerged as a flexible alternative (Farmer and Foley,, 2009), allowing for heterogeneous, adaptive agents whose interactions generate emergent macroeconomic patterns. Within monetary policy research, ABMs have been used to model expectation formation and policy transmission in bottom-up settings (Delli Gatti et al.,, 2011). Yet two key limitations remain. First, behavioral rules are typically imposed ex ante as stylized heuristics, rather than learned or emergent (Horton,, 2023). Second, conventional ABMs cannot interpret qualitative information directly, despite the fact that monetary policy signals are largely communicated in natural language. This omission bypasses an essential channel of transmission. Our framework addresses both limitations by enabling agents to process and act upon policy-relevant textual information.
@@ 3 Methodology and Data
Our methodology proceeds in three main steps: (i) construction of a structured database of ECB communication events and corresponding market reactions; (ii) development of LLM-based synthetic agents embedded in an agent-based simulation framework; and (iii) evaluation of the correspondence between synthetic measures of belief heterogeneity and actual market outcomes. We focus on ECB press conferences, which constitute the primary vehicle for conveying monetary policy decisions in the Euro area. Following the literature on high-frequency identification of monetary policy shocks (Gurkaynak et al.,, 2005; Altavilla et al.,, 2019), we define event dates as the days of regularly scheduled ECB Governing Council monetary policy meetings. For each event, we collect the full transcript of the press conference as published on the ECB website as well as the associated OIS post-conference volatility for the short (3-months), medium (2-years), and long (10-years) term. [^3] Following the monetary policy communication literature, we use OIS rates as they reflect pure interest rate expectations without the credit risk component present in government bond yields (Gurkaynak et al.,, 2005). Our final sample comprises 283 ECB conferences between June 1998 and April 2025.
@@ 3.2 Prompting Strategies
The baseline approach, where the agent receives only the ECB press conference transcript. This approach tests whether LLMs possess intrinsic capabilities to decode monetary policy communication without explicit calibration.
@@
Augments the baseline by including the transcripts of the three most recent ECB press conferences and the associated historical changes in market-based disagreement, providing a short memory of past communication–market interactions.This design tests whether contextual examples improve forecasting accuracy, following established few-shot learning principles. [^6] Few-shot prompting has demonstrated superior performance over zero-shot approaches across multiple economic forecasting applications, particularly when fine-tuning is computationally infeasible (Hansen et al.,, 2024).
@@ 4 Results
This section evaluates the performance of our three prompting strategies in replicating market disagreement patterns following ECB press conferences. We examine the zero-shot baseline (Section 4.1), the few-shot approach with historical context (Section 4.2), and the iterative LLM-as-a-Judge framework (Section 4.3), focusing on correlations between synthetic disagreement and actual market volatility across different yield curve tenors.
@@ 4.1 Zero-Shot Prompt
Note: Each data point reflects the output of an LLM model prompted in a zero-shot setting using Gemini 2.5-Flash, with a temperature parameter of 1. Forecast directions are aggregated across agents. The simulation uses a sample of 283 ECB press conference (June 1998-April 2025) transcripts and the baseline prompt (Figure 2).
@@
Note: Each data point reflects the output of an LLM model prompted in a zero-shot setting using Gemini 2.5-Flash, with a temperature parameter of 1. Dispersion is measured as the cross-sectional standard deviation of rate predictions. The simulation uses a sample of 283 ECB press conferences (June 1998-April 2025) and the baseline prompt (Figure 2). Highlighted points represent observations exceeding the 95th percentile of dispersion for each tenor.
@@
Figure 5 shows the cross-sectional standard deviation of forecasts, revealing distinct phases that align closely with monetary policy regimes and market stress episodes. Dispersion contracts systematically during periods of unambiguous policy stance, such as the post-crisis easing phase and much of the zero lower bound (ZLB) era, when forward guidance provided clear anchoring. In this period, the 3-month dispersion falls below 5 basis points, with longer tenors, instead, clustering around 10-15 basis points. Conversely, it spikes dramatically at policy turning points and during crisis episodes, touching sharp peaks during the 2008-2009 financial crisis (with the 2 and 10-year dispersion exceeding 40 basis points), the 2011-2012 sovereign debt crisis, and most recently during the 2022-2023 monetary policy tightening cycle, where disagreement reached unprecedented levels (roughly 60 basis points for the 2 and 10-year maturity). In the latter instance, disagreement arises less from directional confusion and more from uncertainty over policy speed, magnitude, and terminal endpoints. This result is consistent with the ECB 2025 Monetary Policy Strategy Review findings (ECB,, 2025). The report shows that, once forward guidance was abandoned, the sensitivity of OIS rates to macroeconomic news increased significantly, hence amplifying excess volatility. Moreover, the early phase of the 2022 hiking cycle was marked by historically large monetary policy shocks, including the unexpected 50 basis point hike in July 2022 and the largest hawkish surprise since the global financial crisis in December 2022, reflecting the challenge for the ECB of aligning communication with policy moves. By contrast, following the introduction of the three-element reaction function framework in March 2023, the subsequent phase of the tightening cycle exhibits markedly subdued volatility to policy announcements. These patterns mirror survey-based and market-implied measures of disagreement, where even directionally well-telegraphed policy paths can generate substantial heterogeneity in rate forecasts. [^15] See Gilchrist et al. (2014) and Bayer et al. (2022) for a thorough discussion on forecast disagreement and monetary policy uncertainty. [^16] At the end of 1998, volatility exhibits another peak for the 3-month tenor. In this period, the ECB avoided signaling its initial policy rate for the monetary union, deferring decisions to December. While acknowledging global risks from the Asian crisis, Russian default, and LTCM collapse, President Duisenberg reaffirmed the commitment to price stability, contrasting with the Fed’s three rate cuts between September and November. This ambiguity left markets uncertain and drove volatility in short-term instruments, mainly 1- to 3-month FRAs and DEM-based futures, since euro derivatives like Euribor futures and OIS did not yet exist (ECB,, 1999). The cross-tenor hierarchy remains remarkably consistent across all episodes, with 10-year uncertainty slightly larger than 2-year uncertainty, and both larger than the 3-month dispersion by 2-3 times, reflecting the increasing complexity of information processing and projection uncertainty at longer time horizons.
@@
Note: Correlation between the output of an LLM model prompted in a zero-shot setting using Gemini 2.5-Flash, with a temperature parameter of 1, and a market-based measure of monetary policy uncertainty from Collodel and Kunzmann, (2025). The simulation uses a sample of 283 ECB press conferences (June 1998-April 2025) and the baseline prompt (Figure 2). Rolling correlation is calculated using a 12-month window.